OpenStudy (anonymous):

Can you calculate this to the next step ? \[\frac{(2n+2)(2n+1)(2n)!}{(2n+2)2^{n}n!} = ? \]

OpenStudy (anonymous):

\[\frac{(2n+2)(2n+1)(2n)!}{(2n+2)2^{n}n!} =\frac{(2n+1)(2n)!}{2\times 4\times 6\times...\times2n} \]make sense till here?

OpenStudy (anonymous):

2x3x6x... this part not..

OpenStudy (anonymous):

actually \[(2n+2)2^{n}n! = ?\] would help me too..

OpenStudy (anonymous):

\[\ \ \ n!=1\times2\times3\times...\times n \]\[\ \ \ 2^n=2\times2\times2\times...\times 2 \]\[2^n n!=2\times4\times6\times...\times2n\]

OpenStudy (anonymous):

ok..

OpenStudy (anonymous):

\[\frac{(2n+2)(2n+1)(2n)!}{(2n+2)2^{n}n!} =\frac{(2n+1)(2n)!}{2\times 4\times 6\times...\times2n}=(2n+1)\times(2n-1)\times...\times5\times3\times1\]

OpenStudy (anonymous):

is it not possible to show it like that for example \[(2n+2)2^{n}n! = 4n^{n} + 4^{n} (n!)\] and i dont know if its right.. ?

OpenStudy (anonymous):

its impossible i think

OpenStudy (anonymous):

ok thanks anyway

OpenStudy (anonymous):

welcome :)